OpenStudy (darkbluechocobo):

Help with equation. Mike has created a function explaining the productivity of his company. He knows his employees produce 80 bicycle parts per hour. He also inventoried his stock and knows that he currently has 90 parts ready to be shipped. What function represents the total bicycles parts (b) in terms of the number of hours (h) worked?

OpenStudy (darkbluechocobo):

(h) = 90h + 80 b) = 80b + 90 (h) = 80h + 90 (b) = 90b + 80

OpenStudy (anonymous):

ok so for every hour of work you have 80 part which is represented with 80h and you already have 90 parts. you would represent it like this 80h+90 hope this helped!

OpenStudy (darkbluechocobo):

so since its 80 parts per hour it would be h? and since he already has 90 to be shipped that would be the second part?

OpenStudy (anonymous):

well it should look like this \[(B)=80(h)+90\]

OpenStudy (anonymous):

sorry for the confusion let me restate what i said properly... B is the total parts you have. You make 80 parts per hour 80(h) and have 90 parts already

OpenStudy (darkbluechocobo):

Okay :3 I believe i understand just to make sure I have another one which has the same concept

OpenStudy (anonymous):

alright let me see

OpenStudy (darkbluechocobo):

Mike has created a function explaining the productivity of his company. He knows his employees produce 40 bicycle parts per hour. He also inventoried his stock and knows that he currently has 500 parts ready to be shipped. What function represents the total bicycles parts (b) in terms of the number of hours (h) worked? (h) = 40h + 500 (b) = 40b + 500 (h) = 500h + 40 (b) = 500b + 40 So wouldnt this be A?

OpenStudy (darkbluechocobo):

Using the same stuff from the other equation?

OpenStudy (anonymous):

yes the question are the same except for the numbers.

OpenStudy (anonymous):

is the same*

OpenStudy (anonymous):

show me how you would represent the equation

OpenStudy (darkbluechocobo):

I would use B=40h since you make forty parts an hour and since you have 500 already the equation would be B=40h+500

OpenStudy (anonymous):

correct! you got it!

OpenStudy (darkbluechocobo):

yay :3

OpenStudy (anonymous):

can you handle this now on your own?

OpenStudy (darkbluechocobo):

yus equations liek that seem easy now :p

OpenStudy (anonymous):

great! close the question so it doesn't keep popping up and have a good day!